OpenStudy (anonymous):

8x^2-64y^2-64=0 write the standard form of this equation

OpenStudy (anonymous):

8*x^2-64*y^2-64=0?

OpenStudy (anonymous):

so 0+64/(8/8)/(2/2) = 64 i think

OpenStudy (anonymous):

idk if that helped

OpenStudy (anonymous):

tbh idk what I'm doing lol

Nnesha (nnesha):

first write in this form Ax^2 +By^2 = C

Nnesha (nnesha):

what kind of equation is that hyperbola ? ellipse ? circle ? Parabola ?